OpenStudy (bloomlocke367):

I need a quick refresher, how do you use long division to divide polynomials?

OpenStudy (irishboy123):

post an example :-)

OpenStudy (irishboy123):

then we can do

OpenStudy (bloomlocke367):

divide f(x) by d(x). \(f(x)=x^2-2x+4\) \(d(x)=x-7\)

OpenStudy (bloomlocke367):

@IrishBoy123

OpenStudy (irishboy123):